Solve the equation (f compose g)(x) = 0. (Enter your answers as a comma-separated list.)
f(x) = x2 − x − 6, g(x) = 2x − 4
I plugged g(x) into f(x) and ended up with 2x^2-9x+3 and used the quadratic to find the solutions, but I wasn't getting the correct answer.

Answer by stanbon(75887)

Solve the equation (f o g)(x) = 0. (Enter your answers as a comma-separated list.)
f(x) = x^2 − x − 6, g(x) = 2x − 4
I plugged g(x) into f(x) and ended up with 2x^2-9x+3 and used the quadratic to find the solutions, but I wasn't getting the correct answer.
-----------------
fog(x) = f[g(x)] = f[2x-4] = (2x-4)^2 - (2x-4) - 6
----
= 4x^2 -16x + 16 - 2x + 4 - 6
----
= 4x^2 -18x + 14
----
Solve:: 2x^2 - 9x + 7 = 0
---
(x-1)(2x-7) = 0
----
Solutions:: x = 1 or x = 7/2
---------------
Cheers,
Stan H.
--------------------